### v3

@@ 1 Introduction
Figure 1: Comparisons with other advanced direct-trained SNNs using ResNet-based architectures on CIFAR100 and ImageNet-1k. Our models maintain the same level of model size as our baselines.
Spiking neural networks (SNNs) are considered the third generation of neural networks due to their emulation of the human brain’s information-processing mechanism [33]. The essential uniqueness of SNN is its activation unit, i.e., spiking neuron, which transfers floating-point values into binary spikes, thus making the matrix operations, e.g., convolution, conducted multiplication-free [15]. Similar to binary neural networks (BNN), binary feature maps (spikes) would reduce spatial representation capacity, making processing information-dense data challenging and resulting in lower model accuracy [57, 35, 56].
In this context, multi-bit spiking neural networks have recently arisen as a promising approach to improving such inadequate informational scope [18, 49, 50, 48]. As the name suggests, multi-bit SNNs introduce more bits to represent the original binary spike and devote more addition operations to process the matrix multiplication. Prior arts successfully incorporated this mechanism to improve accuracy, while most neglected the actual memory and computation overhead as pointed out by [39]. For instance, a layer with less representation ability but assigned with higher bit widths would incur memory and computation waste and interfere with the model inference [12, 6]. It necessarily comes to the question: Can we enhance SNN’s performance while keeping memory and computation costs at minimal levels?
@@
In this paper, we investigate the adaptive bit allocation of quantized SNNs as the answer. Specifically, we parametrize the three fundamental elements of memory and computation: temporal lengths, bit widths of weights, and bit widths of spikes, making them learnable and able to be optimized end-to-end. To regulate the allocation process of temporal lengths and bit widths during training and ensure they converge to the targets correctly, we explicitly set temporal length and bit width bounds and implicitly design the regulating loss. By implementing these strategies, we can effectively assign different temporal lengths and bit widths to different layers, enabling a fine-grained allocation of memory and computation resources for SNN. Thus, the accuracy and efficiency of the inference can be improved. Moreover, unlike the memory and computation measurement of most prior arts [56, 18, 50, 52], we incorporate the pragmatic concept "Bit Budget" and the associated S-ACE, which are proposed by [39], to estimate the accurate model memory and computation overheads.
@@ 2 Related work
Based on the idea that SNNs could be optimized end-to-end through back-propagation [37], Bothe et al. [5] first used the surrogate gradient (SG) to solve the non-differentiability of SNNs. Wu et al. [45] further compared and analyzed the impacts of the shapes of SG functions on model performance. Afterward, many practical techniques and architectural designs were developed to strengthen the direct-trained SNN performance, such as tdBN [55], SEW block [15], spiking self-attention (SSA) block [56], TET [11], etc. In this paper, we adopt the two main-stream direct-trained SEW and SSA to estimate the effectiveness of our proposed methods.
@@
Wu et al. [47] were the first to replace binary spikes with full-precision analog spikes to improve the model accuracy. Guo et al. added the spike amplitude coefficient [19] and negative spike [18] to realize ternary (2-bit) spikes. You et al. [50] and Xing et al. [50] concurrently introduced ternary spikes into natural language processing (NLP), and both achieved tremendous success. SpikeLLM [49] and Multi-bit SNN [48] were more aggressive. They extended ternary spikes to multiple bits (over 2 bits) to boost the SNN performance while theoretically proving that the characteristic of addition-only computation can still be maintained. Shen et al. [39] combined multi-bit SNNs with quantization and presented a rational and practical approach called "Bit Budget" to estimating computing overheads. Different from the above multi-bit SNNs, our work focuses on fine-grained bit allocations, with which multi-bit SNNs can substantially minimize the memory and computation and maximize the model accuracy.
@@
Mixed-precision quantization is based on the insight that different layers in artificial neural networks (ANN) require different degrees of representation ability. Wang et al. [43] and Wu et al. [44] were the pioneers that fostered the search-based means to allocate different bit widths with reinforcement learning and differentiable neural architecture search, respectively. While, Dong et al. [12] used Hessian metrics to represent the layer importance and accordingly assigned bit widths. Recently, optimization-based mixed-precision was developed to learn the bit width during training [42, 54, 24, 25]. Compared with the above methods, which focus solely on ANN mixed-precision, our work is among the first to target SNNs, addressing the additional temporal dimension and mitigating step-size mismatch.
@@ 3.1 Multi-bit spiking neuron
Rooted from burst encoding [30], the multi-bit spiking neuron generalizes Equation 2 to allow multiple spikes at the same time-step. The formulation transforms into [39, 48, 49]:
@@
For further improvement of SNN’s efficiency, prior work has employed weight quantization to compress the model size, e.g., [39]. A classic quantization formula is given below:
@@
Shen et al. [39] propose the "Bit Budget" paradigm to fairly and practically estimate the substantial computation and memory. Following is the definition:
@@
To measure the computational consumption of fully ideal neuromorphic hardware. Shen et al. [39] also propose the NS-ACE metric, which is calculated by:
$\begin{split}\text{NS-ACE}&=\sum_{w\in W,s\in S}fr_{s}\cdot n_{w,s}\cdot\text{BB}\\ &\approx\bar{fr}_{s}\cdot\sum_{w\in W,s\in S}n_{w,s}\cdot\text{BB}\\ &=\bar{fr}_{s}\cdot\text{S-ACE}.\end{split}$ (9)
Here, $fr_{s}$ ($\bar{fr}_{s}$ ) is the (average) firing rate of neurons, which is always less than $1$.
@@ 4.1 Parametrization
We can observe that Equation 5 is essentially a flooring quantization process of $v^{t}_{l}$, where $S_{out,l}^{t}$ is the quantized integer, $V_{th}$ is the quantization step size, and the quantized spike value is $\alpha\cdot S_{out}^{t}$. Therefore, SNN’s feature map size is controlled by $B_{s}$ and $T$. Similarly, the weight parameter size can also be controlled by the weight bit width $B_{w}$ in Equation 6. Thus, the whole model’s computation and memory overhead is controlled by $B_{s}$, $T$, and $B_{w}$, which is why the "Bit budget" metric [39] is valid and matters. However, previous SNN quantization [39] did not take into account a fine-grained allocation mechanism for these pivotal elements ($B_{s}$, $T$, and $B_{w}$), leading to suboptimal model efficiency.
Based on the insight that a layer with less representation ability but assigned with higher bit widths would incur memory and computation waste and interfere with the model inference [43, 6], we propose to parametrize the three pivotal elements layer-wise, making them learnable and can be optimized through back-propagation. Thus, a fine-grained allocation strategy would be feasible. Following are the parametrization formulas, including:
@@ 4.2 Formulation of the refined parametric spiking neuron
Firstly, since the temporal length is made parametric and can be different among different layers during training, two issues arise: (1) the gradient requirement, i.e., $T_{l}$ needs to get involved in the forward pass; (2) the temporal mismatch between layers. For example, the previous layer yields a spike train with $T=2$, while the post layer has four temporal bits, i.e., $T=4$. Two time-steps of data remain unfilled. To solve these issues, we exploit the temporal flexibility [51], directly squeezing the spike train through an averaging operation, as shown in Figure 2(a) and Equation 13. Notably, the extra computation brought by such temporal squeezing is negligible because it only makes up less than 0.1% of the total model computation. In addition, the spiking self-attention[56] layer requires the three spike feature maps ($Q$, $K$, and $V$) to perform matrix multiplication, necessitating compatible feature map dimensions. Therefore, we enforce shared $T_{l}$ among the corresponding spiking neurons to ensure the intra-layer temporal matching, as shown in Figure 2(a).
@@
Thirdly, the spike amplitude coefficient $\alpha$ is usually defined differently among prior arts [39, 18, 48]. Here, we keep consistent with quantization’s perspective, and directly unify $V_{th}$ and $\alpha$ as identical, layer-wise, learnable quantization step size, as shown in Equations 13- 15.
@@
where, in our practice, $\tau$ is suggested to be 1, the spike amplitude coefficient $\alpha$ is replaced by the learnable membrane-potential threshold $V^{1}_{th,l}$, and threshold shift $V^{2}_{th,l}$ equates to $0.5sign(v^{t}_{l}/V^{1}_{th,l})$. Beyond this, to further increase the adaptivity [53, 50], we apply the temporal-wise sharing technique to $V_{th,l}^{1}$ and $B_{s,l}$. Following [53] and [50], we make $V_{th,l}^{1}$ and $B_{s,l}$ shared temporally along the temporal dimension, as shown in Equation 15. Specifically, $V_{th,l}^{1}\Rightarrow V_{th,l}^{1,t}\in\{V_{th,l}^{1,1},...,V_{th,l}^{1,t},...,V_{th,l}^{1,T_{bound}}\}.$ Similarly, $B_{s,l}\Rightarrow B_{s,l}^{t}\in\{B_{s,l}^{1},...,B_{s,l}^{t},...,B_{s,l}^{T_{bound}}\}.$
As suggested in [19, 18, 32], negative spikes should be enabled to improve the representation ability of SNNs. Therefore, we expand Equation 15 to the bidirectional domain, thus supporting negative spikes. Equation 15 would turn into:
@@ 4.3.1 Gradient calculations
We solve non-differentiation via straight-through estimator (STE) [3] and chain rule.
@@
Here, $j$ is the index of the feature map element. The minimum quantization integer $q_{min}=0$ and the maximum integer $q_{max}=2^{B_{s,l}^{t}}-1$. While, $g_{s,scale}$, equal to $\frac{1}{\sqrt{\sum_{j}1\cdot q_{max}}}$ and meant to avoid gradient explosion [13], is the gradient scaling factor. $L_{task}$ represents the task loss signal.
@@
3) For $V^{1}_{th,l}$ and $S^{l}_{q}$, the derivation of gradients is similar to prior arts [13] and trivial, so we defer it to the appendix.
@@ 4.4.1 Quantization step size mismatch issue
With the conclusion I) and II) combined, the step-size mismatch issue can be proven harmful in direct training, where bit-width mutation constantly occurs between forward passes as shown in Figure 3. As the model goes deeper, quantization error accumulates and seriously impairs the model output [38]. The step size mismatch influence could be more severe in SNNs because the invalid step size will be consecutively used along the temporal dimension in Equation 5. Therefore, the quantization error will accumulate temporally, causing serious inference offsets. For instance, if $X=\{x_{1},x_{2},...,x_{T}\}$, then the possibility of a increased quantization error $(X-X_{q})^{2}$ would be $1-[1-P(x>\frac{2^{b^{\prime}}-1}{2^{b}-1}3\sigma)]^{T}$, which increases as $T$ rises.
@@ 4.4.2 Step-size renewal mechanism
Table 1: Comparisons with existing works on CIFAR. W/S/T denotes the averaged weight/spike/temporal bit width. U-Quant., following [39], is the plain uniform quantization on SNN, which uses fixed uniform bit widths. Our models are initial to W/S/T=4/4/2.
@@
Method Architecture W/S/T Bit Budget S-ACE (G) NS-ACE(G) Size (MB) Top-1 Hybrid training [36] ResNet34 16/1/250 4000 13608 - 43.56 61.48 Spiking ResNet [23] ResNet50 16/1/350 5600 18952 - 51.12 72.75 tdBN [55] ResNet34 16/1/6 96 340.39 - 43.56 63.72 Dspike [29] ResNet34 16/1/6 96 340.39 - 43.56 68.19 MS-ResNet [22] ResNet34 16/1/6 96 340.39 - 43.56 69.42 Attention-SNN [52] ResNet34 16/1/1 16 56.73 - 43.56 69.15 TET [11] Sew-ResNet34 16/1/4 64 226.93 - 43.56 68.00 Ternary Spike [18] ResNet34 16/2/4 128 435.86 - 43.56 70.74 PLIF+SEW [15] SEW-ResNet34 16/1/4 64 226.93 63.14 43.56 67.04 SEW-ResNet50 16/1/4 64 215.62 69.03 51.12 67.78 Quant. SNN [39] SEW-ResNet34 8/8/1 64 226.93 56.42 21.78 70.13 Ours-221 SEW-ResNet34 2.58/2.18/1.0 5.61 30.74 7.77 7.03 70.57 Ours-241 2.62/3.96/1.0 10.39 54.69 16.99 7.14 72.02 Ours-441 3.93/3.92/1.0 15.4 77.78 24.30 10.69 72.82

### v4

@@ 1 Introduction
Figure 1: Comparisons with other advanced direct-trained SNNs using ResNet-based architectures on CIFAR100 and ImageNet-1k. Our models maintain the same level of model size as the baselines.
Spiking neural networks (SNNs) are considered the third generation of neural networks due to their emulation of the human brain’s information-processing mechanism [maass1997networks]. The essential uniqueness of SNN is its activation unit, i.e., spiking neuron, which transfers floating-point values into binary spikes, thus making the matrix operations, e.g., convolution, conducted multiplication-free [fang2021deep]. Similar to binary neural networks (BNN), binary feature maps (spikes) would reduce spatial representation capacity, making processing information-dense data challenging and resulting in lower model accuracy [zhu2019binary, meng2022training, zhou2022spikformer].
In this context, multi-bit spiking neural networks have recently arisen as a promising approach to improving such inadequate informational scope [guo2024ternary, xing2024spikellm, xing2024spikelm, xiao2024multi]. As the name suggests, multi-bit SNNs introduce more bits to represent the original binary spike and devote more addition operations to process the matrix multiplication. Prior arts successfully incorporated this mechanism to improve accuracy, while most neglected the actual memory and computation overhead as pointed out by [shen2024conventional]. For instance, a layer with less representation ability but assigned with higher bit widths would incur memory and computation waste and interfere with the model inference [dong2019hawq, Chen_2021_ICCV]. It necessarily comes to the question: Can we enhance SNN’s performance while keeping memory and computation costs at minimal levels?
@@
In this paper, we investigate the adaptive bit allocation of quantized SNNs as the answer. Specifically, we parametrize the three fundamental elements of memory and computation: temporal lengths, bit widths of weights, and bit widths of spikes, making them learnable and able to be optimized end-to-end. To regulate the allocation process of temporal lengths and bit widths during training and ensure they converge to the targets correctly, we explicitly set temporal length and bit width bounds and implicitly design the regulating loss. By implementing these strategies, we can effectively assign different temporal lengths and bit widths to different layers, enabling a fine-grained allocation of memory and computation resources for SNN. Thus, the accuracy and efficiency of the inference can be improved. Moreover, unlike the memory and computation measurement of most prior arts [zhou2022spikformer, guo2024ternary, xing2024spikelm, yao2023attention], we incorporate the pragmatic concept "Bit Budget" and the associated S-ACE, which are proposed by [shen2024conventional], to estimate the accurate model memory and computation overheads.
@@ 2 Related work
Based on the idea that SNNs could be optimized end-to-end through back-propagation [rumelhart1986learning], Bothe et al. [bohte2000spikeprop] first used the surrogate gradient (SG) to solve the non-differentiability of SNNs. Wu et al. [wu2018spatio] further compared and analyzed the impacts of the shapes of SG functions on model performance. Afterward, many practical techniques and architectural designs were developed to strengthen the direct-trained SNN performance, such as tdBN [zheng2021going], SEW block [fang2021deep], spiking self-attention (SSA) block [zhou2022spikformer], TET [deng2021temporal], etc. In this paper, we adopt the two main-stream direct-trained SEW and SSA to estimate the effectiveness of our proposed methods.
@@
Wu et al. [wu2021liaf] were the first to replace binary spikes with full-precision analog spikes to improve the model accuracy. Guo et al. added the spike amplitude coefficient [guo2022real] and negative spike [guo2024ternary] to realize ternary (2-bit) spikes. You et al. [xing2024spikelm] and Xing et al. [xing2024spikelm] concurrently introduced ternary spikes into natural language processing (NLP), and both achieved tremendous success. SpikeLLM [xing2024spikellm] and Multi-bit SNN [xiao2024multi] were more aggressive. They extended ternary spikes to multiple bits (over 2 bits) to boost the SNN performance while theoretically proving that the characteristic of addition-only computation can still be maintained. Shen et al. [shen2024conventional] combined multi-bit SNNs with quantization and presented a rational and practical approach called "Bit Budget" to estimating computing overheads. Different from the above multi-bit SNNs, our work focuses on fine-grained bit allocations, with which multi-bit SNNs can substantially minimize the memory and computation and maximize the model accuracy.
@@
Mixed-precision quantization is based on the insight that different layers in artificial neural networks (ANN) require different degrees of representation ability. Wang et al. [wang2019haq] and Wu et al. [wu2018mixed] were the pioneers that fostered the search-based means to allocate different bit widths with reinforcement learning and differentiable neural architecture search, respectively. While, Dong et al. [dong2019hawq] used Hessian metrics to represent the layer importance and accordingly assigned bit widths. Recently, optimization-based mixed-precision was developed to learn the bit width during training [uhlich2019mixed, zhang2021differentiable, huang2022sdq, kim2024metamix]. Compared with the above methods, which focus solely on ANN mixed-precision, our work is among the first to target SNNs, addressing the additional temporal dimension and mitigating step-size mismatch.
@@ 3.1 Multi-bit spiking neuron
Rooted from burst encoding [liefficient], the multi-bit spiking neuron generalizes Equation 2 to allow multiple spikes at the same time-step. The formulation transforms into [shen2024conventional, xiao2024multi, xing2024spikellm]:
@@
For further improvement of SNN’s efficiency, prior work has employed weight quantization to compress the model size, e.g., [shen2024conventional]. A classic quantization formula is given below:
@@
Shen et al. [shen2024conventional] propose the "Bit Budget" paradigm to fairly and practically estimate the substantial computation and memory. Following is the definition:
@@
To measure the computational consumption of fully ideal neuromorphic hardware, Shen et al. [shen2024conventional] also propose the NS-ACE metric to ignore the computation cost of non-spike data, which is calculated by:
$\begin{split}\text{NS-ACE}&=\sum_{w\in W,s\in S}fr_{s}\cdot n_{w,s}\cdot\text{BB}\\ &\approx\overline{fr}_{s}\cdot\sum_{w\in W,s\in S}n_{w,s}\cdot\text{BB}\\ &=\overline{fr}_{s}\cdot\text{S-ACE}.\end{split}$ (9)
Here, $fr_{s}$ ($\overline{fr}_{s}$ ) is the (average) firing rate of neurons, which is always less than $1$.
@@ 4.1 Parametrization
We can observe that Equation 5 is essentially a flooring quantization process of $v^{t}_{l}$, where $S_{out,l}^{t}$ is the quantized integer, $V_{th}$ is the quantization step size, and the quantized spike value is $\alpha\cdot S_{out}^{t}$. Therefore, SNN’s feature map size is controlled by $B_{s}$ and $T$. Similarly, the weight parameter size can also be controlled by the weight bit width $B_{w}$ in Equation 6. Thus, the whole model’s computation and memory overhead is controlled by $B_{s}$, $T$, and $B_{w}$, which is why the "Bit budget" metric [shen2024conventional] is valid and matters. However, previous SNN quantization [shen2024conventional] did not take into account a fine-grained allocation mechanism for these pivotal elements ($B_{s}$, $T$, and $B_{w}$), leading to suboptimal model efficiency.
Based on the insight that a layer with less representation ability but assigned with higher bit widths would incur memory and computation waste and interfere with the model inference [wang2019haq, Chen_2021_ICCV], we propose to parametrize the three pivotal elements layer-wise, making them learnable and can be optimized through back-propagation. Thus, a fine-grained allocation strategy would be feasible. Following are the parametrization formulas, including:
@@ 4.2 Formulation of the refined parametric spiking neuron
Firstly, since the temporal length is made parametric and can be different among different layers during training, two issues arise: (1) the gradient requirement, i.e., $T_{l}$ needs to get involved in the forward pass; (2) the temporal mismatch between layers. For example, the previous layer yields a spike train with $T=2$, while the post layer has four temporal bits, i.e., $T=4$. Two time-steps of data remain unfilled. To solve these issues, we exploit the temporal flexibility [yao2021temporal], directly squeezing the spike train through an averaging operation, as shown in Figure 2(a) and Equation 13. Notably, the extra computation brought by such temporal squeezing is negligible because it only makes up less than 0.1% of the total model computation. In addition, the spiking self-attention[zhou2022spikformer] layer requires the three spike feature maps ($Q$, $K$, and $V$) to perform matrix multiplication, necessitating compatible feature map dimensions. Therefore, we enforce shared $T_{l}$ among the corresponding spiking neurons to ensure the intra-layer temporal matching, as shown in Figure 2(a).
@@
Thirdly, the spike amplitude coefficient $\alpha$ is usually defined differently among prior arts [shen2024conventional, guo2024ternary, xiao2024multi]. Here, we keep consistent with quantization’s perspective, and directly unify $V_{th}$ and $\alpha$ as identical, layer-wise, learnable quantization step size, as shown in Equations 13- 15.
@@
where, in our practice, $\tau$ is suggested to be 1, the spike amplitude coefficient $\alpha$ is replaced by the learnable membrane-potential threshold $V^{1}_{th,l}$, and threshold shift $V^{2}_{th,l}$ equates to $0.5sign(v^{t}_{l}/V^{1}_{th,l})$. Beyond this, to further increase the adaptivity [yao2022glif, xing2024spikelm], we apply the temporal-wise sharing technique to $V_{th,l}^{1}$ and $B_{s,l}$. Following [yao2022glif] and [xing2024spikelm], we make $V_{th,l}^{1}$ and $B_{s,l}$ shared temporally along the temporal dimension, as shown in Equation 15. Specifically, $V_{th,l}^{1}\Rightarrow V_{th,l}^{1,t}\in\{V_{th,l}^{1,1},...,V_{th,l}^{1,t},...,V_{th,l}^{1,T_{bound}}\}.$ Similarly, $B_{s,l}\Rightarrow B_{s,l}^{t}\in\{B_{s,l}^{1},...,B_{s,l}^{t},...,B_{s,l}^{T_{bound}}\}.$
As suggested in [guo2022real, guo2024ternary, liu2022spikeconverter], negative spikes should be enabled to improve the representation ability of SNNs. Therefore, we expand Equation 15 to the bidirectional domain, thus supporting negative spikes. Equation 15 would turn into:
@@ 4.3.1 Gradient calculations
We solve non-differentiation via straight-through estimator (STE) [bengio2013estimating] and chain rule.
@@
Here, $j$ is the index of the feature map element. The minimum quantization integer $q_{min}=0$ and the maximum integer $q_{max}=2^{B_{s,l}^{t}}-1$. While, $g_{s,scale}$, equal to $\frac{1}{\sqrt{\sum_{j}1\cdot q_{max}}}$ and meant to avoid gradient explosion [esser2020learned], is the gradient scaling factor. $L_{task}$ represents the task loss signal.
@@
3) For $V^{1}_{th,l}$ and $S^{l}_{q}$, the derivation of gradients is similar to prior arts [esser2020learned] and trivial, so we defer it to the appendix.
@@ 4.4.1 Quantization step size mismatch issue
With the conclusion I) and II) combined, the step-size mismatch issue can be proven harmful in direct training, where bit-width mutation constantly occurs between forward passes as shown in Figure 3. As the model goes deeper, quantization error accumulates and seriously impairs the model output [sengupta2019going]. The step size mismatch influence could be more severe in SNNs because the invalid step size will be consecutively used along the temporal dimension in Equation 5. Therefore, the quantization error will accumulate temporally, causing serious inference offsets. For instance, if $X=\{x_{1},x_{2},...,x_{T}\}$, then the possibility of a increased quantization error $(X-X_{q})^{2}$ would be $1-[1-P(x>\frac{2^{b^{\prime}}-1}{2^{b}-1}3\sigma)]^{T}$, which increases as $T$ rises.
@@ 4.4.2 Step-size renewal mechanism
Table 1: Comparisons with existing works on CIFAR. W/S/T denotes the averaged weight/spike/temporal bit width. U-Quant., following [shen2024conventional], is the plain uniform quantization on SNN, which uses fixed uniform bit widths. Our models are initial to W/S/T=4/4/2.
@@
Method Architecture W/S/T Bit Budget S-ACE (G) NS-ACE(G) Size (MB) Top-1 Hybrid training [rathi2020enabling] ResNet34 16/1/250 4000 13608 - 43.56 61.48 Spiking ResNet [hu2021spiking] ResNet50 16/1/350 5600 18952 - 51.12 72.75 tdBN [zheng2021going] ResNet34 16/1/6 96 340.39 - 43.56 63.72 Dspike [li2021differentiable] ResNet34 16/1/6 96 340.39 - 43.56 68.19 MS-ResNet [hu2024advancing] ResNet34 16/1/6 96 340.39 - 43.56 69.42 Attention-SNN [yao2023attention] ResNet34 16/1/1 16 56.73 - 43.56 69.15 TET [deng2021temporal] Sew-ResNet34 16/1/4 64 226.93 - 43.56 68.00 Ternary Spike [guo2024ternary] ResNet34 16/2/4 128 435.86 - 43.56 70.74 PLIF+SEW [fang2021deep] SEW-ResNet34 16/1/4 64 226.93 63.14 43.56 67.04 SEW-ResNet50 16/1/4 64 215.62 69.03 51.12 67.78 Quant. SNN [shen2024conventional] SEW-ResNet34 8/8/1 64 226.93 56.42 21.78 70.13 Ours-221 SEW-ResNet34 2.58/2.18/1.0 5.61 30.74 7.77 7.03 70.57 Ours-241 2.62/3.96/1.0 10.39 54.69 16.99 7.14 72.02 Ours-441 3.93/3.92/1.0 15.4 77.78 24.30 10.69 72.82
5 Experiments
In this section, we first compare our methods with the existing works using ResNet on CIFAR [krizhevsky2009learning] and ImageNet-1k [deng2009imagenet], then conduct ablation studies to verify different aspects of our methods, including (1) each proposed technology’s verification, (2) our methods’ extension to spiking self-attention [zhou2022spikformer], (3) mitigating the "temporal constraint" of the prior arts [shen2024conventional] on the event dataset CIFAR10-DVS [li2017cifar10] and DVS-GESTURE [amir2017low], (4) the learned bit allocation’s visualization, and (5) comparing the mixed-precision ANNs. All implementation details are attached in the appendix.
5.1 Comparisons with existing works
5.1.1 CIFAR
We apply our methods to ResNet20 [guo2024ternary] and ResNet18 [li2021differentiable] to compete with the advanced SNNs. As shown in Table 1, we first compare our methods with those using extra learning- or module-based model optimization [li2021differentiable, deng2021temporal, zheng2021going], and achieve superior accuracy using much lower memory (bit budget) and computation (S-ACE). In the case of ResNet-18 with target W/S/T being 1/1/1, our methods can also compress the model sizes to one-tenth of their original dimensions, while achieving at most 2% and 3.74% higher model accuracy on CIFAR-10 and CIFAR-100, respectively. Such a reduction of storage requirements would make SNNs more compatible with on-chip computing architectures [akopyan2015truenorth, mao2024stellar]. Comparing with Ternary Spike ResNet20 [guo2024ternary] that adopts 2-bit bidirectional spike, we can outperform it by at least 1.19% and 5.88% accuracy increase on CIFAR-10 and -100, respectively, while using average 1.32- and 1.58-bit spike and reducing 91.37% and 90.59% model sizes. Then, we compare with Multi-bit ResNet20 [xiao2024multi], whose memory and computation respectively increase to huge 256 bit budgets and 660.48 G S-ACEs. While, our model can achieve not only 1.22% and 3.39% accuracy increase on CIFAR-10 and CIFAR-100 but also two magnitudes of memory and computation savings. Finally, compared with plain uniform quantization (U-quant.), our methods can maximize the model accuracy while minimizing memory and computation requirements. The advantages of our approaches in terms of memory usage, computational efficiency, and storage requirements are consistently demonstrated. Additionally, since we adopt temporal squeezing as shown in Figure 2 and Equation 13, the temporal dimension can be compressed to 1 time-step before operating convolution. That is why S-ACE can further decrease when "W/S/T=222" and more elaborations are in the appendix for better understanding. In conclusion, by comparing with those advanced methods on CIFAR-10 and CIFAR-100, we demonstrate significant improvements in memory efficiency, computational requirements, and storage consumption.
5.1.2 ImageNet
We further apply our methods to SEW-ResNet34 to demonstrate superiority on the up-scaled dataset. As listed in Table 2, compared with those using improved learning techniques or special modules [rathi2020enabling, zheng2021going, li2021differentiable, hu2024advancing], we can achieve magnitudes of memory and computation savings without using any of their tricks. Attention SNN [yao2023attention] has comparable S-ACE (56.73 G) with ours (30.74 G and 54.69 G), but its memory efficiency, model accuracy, and storage requirement lag far behind those of our methods. We can achieve at least 1.42% increased accuracy while reducing 45.81% S-ACE and 83.86% model size. Besides, our models would not require sophisticated attention modules [yao2023attention] as they do, which indicates a more straightforward inference dataflow and a better computing flow mapping from software to hardware. Furthermore, we can save 91.2% memory (5.61 v.s. 64) and 85.83% computation (30.74 G v.s. 217 G) while achieving a higher accuracy (70.57% v.s. 70.13%) when comparing with the quantized SNN [shen2024conventional]. As the bit budget increases, the model accuracy can be advanced to 72.82% with low memory and computation overheads, outperforming full-precision SNNs [guo2024ternary, fang2021deep, deng2021temporal]. Conclusively, scaling our model to the large ImageNet-1k dataset further demonstrates the effectiveness of the proposed methods in improving the efficiency and accuracy of SNNs.